OpenStudy (anonymous):

Prove that sin(x+(pi/4))-sin(x-(pi/4))=sqrt2 cos(x)

OpenStudy (anonymous):

Use the following identity: \[\sin(\alpha\pm\beta)=\sin\alpha\cos\beta\pm\sin\beta\cos\alpha\]

OpenStudy (anonymous):

Where does the sqrt come from?

Directrix (directrix):

If you apply the formulas to the left side: sin(x+(pi/4))-sin(x-(pi/4))=, you'll see that the square root comes up when you crank out the sine of (pi/4). First, you have to apply the formulas. @tbomgard